OpenStudy (anonymous):

The series 1 k(k+1) k = 1 can be rewritten in the form 1 k − 1 k+1 k = 1 using partial fraction decomposition. a) Find the fourth partial sum S4 b) Find the nth partial sum Sn

OpenStudy (anonymous):

Here what we have \[ \sum _{k=1}^n \frac{1}{k (k+1)}=\sum _{k=1}^n \left(\frac{1}{k}-\frac{1}{k+1}\right) \]

OpenStudy (anonymous):

\[ S_4=\sum _{k=1}^4 \left(\frac{1}{k}-\frac{1}{k+1}\right)=\\ 1-\frac 1 2+\\ \frac 1 2 -\frac 13 +\\ \frac 1 3 -\frac 14 +\\ \frac 1 4-\frac 15 \\= 1-\frac 1 5=\frac 4 5 \] By telescopping

OpenStudy (anonymous):

\[ S_n=\sum _{k=1}^n \left(\frac{1}{k}-\frac{1}{k+1}\right)= 1-\frac 1 {n+1} \] All terms geta nnihilated except 1 and\(\large -\frac{ 1} {n+1}\)